# Enhancing Virtual Try-On with Synthetic Pairs and Error-Aware Noise Scheduling

### v1

@@ Abstract
Given an isolated garment image in a canonical product view and a separate image of a person, the virtual try-on task aims to generate a new image of the person wearing the target garment. Prior virtual try-on works face two major challenges in achieving this goal: a) the paired (human, garment) training data has limited availability; b) generating textures on the human that perfectly match that of the prompted garment is difficult, often resulting in distorted text and faded textures. Our work explores ways to tackle these issues through both synthetic data as well as model refinement. We introduce a garment extraction model that generates (human, synthetic garment) pairs from a single image of a clothed individual. The synthetic pairs can then be used to augment the training of virtual try-on. We also propose an Error-Aware Refinement-based Schrödinger Bridge (EARSB) that surgically targets localized generation errors for correcting the output of a base virtual try-on model. To identify likely errors, we propose a weakly-supervised error classifier that localizes regions for refinement, subsequently augmenting the Schrödinger Bridge’s noise schedule with its confidence heatmap. Experiments on VITON-HD and DressCode-Upper demonstrate that our synthetic data augmentation enhances the performance of prior work, while EARSB improves the overall image quality. In user studies, our model is preferred by the users in an average of 59% of cases.
@@ 1 Introduction
Virtual try-on aims to generate a photorealistic image of a target person wearing a prompted product-view garment [23, 40, 37]. It allows users to visualize how garments would fit and appear on their bodies without the need for physical trials. While recent methods have made significant strides in this field [30, 36, 37, 19], noticeable artifacts such as text distortion and faded textures persist in generated images. For example, as illustrated in the second row of Fig. 1, the logo and the text on the t-shirt noticeably fade away in the initial image generated by a prior try-on model [30]. These imperfections stem from two primary challenges in virtual try-on: limited data availability and the complexity of accurate garment texture deformation. Addressing the issues, we propose a two-pronged approach: augmenting training data through cost-effective synthetic data generation, and surgically targeting known generation artifacts using our proposed Error-Aware Refinement-based Schrödinger Bridge (EARSB).
At a minimum, the training data of virtual try-on requires paired (human, product-view garment) images. The product-view garment image is a canonical, front-facing view of the clothing with a clean background. A substantial amount of data is needed to capture the combinatorial space comprising all possible human poses, skin tones, viewing angles, and their respective physical interactions with fabric textures, shapes, letterings, and other material properties. Unfortunately, these images are generally available only on copyright-protected product webpages and therefore not readily available for use. To mitigate this issue, we propose to augment training with synthetic data generated from the easier symmetric human-to-garment task, wherein we train a garment-extraction model to extract a canonical product-view garment image from an image of a clothed person. This will allow us to create synthetic paired training data from unpaired datasets [25, 12, 35]. Our results demonstrate that incorporating the more readily available synthetic training pairs can improve image generation quality in the virtual try-on task.
@@
Two components are necessary to achieve such a pipeline: a classifier for identifying localized generation errors, and a refinement model that can re-synthesize content specifically in these localized regions. We found that an effective Weakly-Supervised Classifier (WSC) can be constructed with just a few hours of manual labeling of generation errors. Another benefit of this approach is that it can be easily tailored for the errors of a specific model that produces images with artifacts. The resulting WSC will produce an error map highlighting low-quality regions. Subsequently, we adopt an Image-to-Image Schrodinger Bridge (I2SB) [24] to learn the refinement of these regions in the generated images. While typical diffusion models map from noise to data, I2SB constructs a Schrödinger Bridge (SB) that allows us to map from data to data, or in our setup, generations with artifacts to ground truth images. In addition, we introduce an adaptive noise schedule to direct the SB process to focus on the localized errors by incorporating the classifier’s prediction error into the noise schedule, which we describe in more detail in Sec. 5.1. In the first row of Fig. 1, our refinement SB model (i.e., EARSB) corrects the distorted text in the initially generated image.
@@ 2 Related Work
The addition of synthetic data is often an effective means of improving downstream task performance when it is difficult to amass real data at the necessary scale. This has been demonstrated in the domains of image generation [18, 31] and image editing [5, 34]. Careful applications can also be used to ameliorate dataset imbalance issues, as shown in [10]. Other works such as [1] use self-synthesized data to provide negative guidance for the diffusion model. Our incorporation of synthetic data in the virtual try-on task tackles a specific sub-problem in the broader image editing domain and is similar in spirit to [5, 34]. Specifically, we aim to synthesize paired training data that satisfies the stringent requirements of virtual try-on paired training data – a canonical product-view garment image paired with an example of it being worn. Images of people in clothing are readily available, but it is difficult to obtain a product-view image of the exact clothing they are wearing. To address this, our work proposes and tackles the human-to-garment, which is roughly symmetric to the virtual try-on task and aims to extract the clothing from a person’s photo and project it to the canonical product view.
@@
3 Virtual Try-On Task Definition
@@ 4 Augmented Training with Synthetic Data
Acquiring high-quality pairs $(x_{0},C)$ at scale is challenging due to copyright and brand protection. On the other hand, acquiring images of just humans ($x_{0}$) at scale is considerably more feasible [25, 12, 35, 22]. This observation motivates our proposal of the human-to-garment task to extract a synthetic canonical view image $\hat{C}$ from $x_{0}$. We can then augment our virtual try-on training with $(x_{0},\hat{C})$ pair, requiring only single human images. In the following, we explain the architecture of our human-to-garment model in Sec. 4.1, how we use it to construct the synthetic dataset in Sec. 4.2, and how the synthetic data is used to augment the virtual try-on training in Sec. 4.3.
@@ 4.3 Augmented Virtual Try-on Training
To further prevent distribution leakage of incorporating synthetic data, we explore two means of limiting the effect of the real-synthetic domain gap: (a) two training stages involving pretraining the try-on model using synthetic pairs, and then finetuning on real pairs [20]; (b) training simultaneously on real and synthetic data, but conditioning the try-on model on a real/synthetic flag, similarly to [17]. We found empirically that the second augmentation performs slightly better than the first (See Sec. 6.1).
@@ 5 Error-Aware Refinement Schrödinger Bridge
Apart from the synthetic data augmentation, our second approach to enhancing existing try-on methods is a refinement pipeline in two steps: identifying the errors in the generated try-on image, referred to as $x_{1}$, and then fitting a targeted model that learns to refine the erroneous locations in $x_{1}$. Given some base model $F_{base}(\bar{x}_{0},C,\phi)\to x_{1}$ where $F_{base}()$ can be any pretrained GAN or diffusion-based approach to virtual try-on, $x_{1}$ closely approximates the true real human image $x_{0}$ with some generation artifacts. To automatically identify the artifacts in $x_{1}$, we construct a weakly-supervised error classifier $\text{WSC}(x_{1},C)\to M$ as in Fig. 3, where $M$ is a confidence map predicting a heatmap for likely generation errors. Then, we perform the final refinement by fitting a Schrödinger bridge based on I2SB [24] that can perform the following mapping: $F_{EARSB}(x_{1},C,M,\phi;\theta)\to x_{0}$.
The approach is weakly inspired by boosting methods in that we wish to fit a targeted refinement model that is trained specifically on the generation errors of an existing model. The refinement goal applies to the general setting where we want to refine a flawed image output, though we focus on virtual try-on for this work.
As in Fig. 3, the training of EARSB includes three steps:
@@
In the following, we start with explaining step 2 as step 1 is simply running prior try-on models.
@@ 5.1 Obtaining the Error Map
We start by obtaining the error map $M$ that highlights the corrupted or incorrect area of the initial image $x_{1}$. A weakly-supervised classifier (WSC) is introduced to predict this error map.
@@
The trained WSC will predict an error map $M$ for the initial image $x_{1}$, which is then used to adjust the noise schedule in the following diffusion process.
@@ 5.2 Error-Map-Reweighted SB Formulation
Our choice of locally scaling the noise schedule is based on several observations. For correctly generated regions in $x_{1}$ we want the model to directly copy pixels over to $x_{0}$. As such, it would be wasteful to train the model to add and remove noise from these regions. On the other hand, erroneous regions in the initial try-on images, specifically the more noticeable ones, include generation errors that share little to no structural similarity to the target. These errors include examples such as deformed limbs, and distorted textures/fabrics, which in many cases may need more noise added to prevent the model from conditioning too strongly on the original pattern.
As such, we construct our refinement model $F_{EARSB}(x_{1},C,M,P;\theta)\to x_{0}$, where the Schrödinger is conditioned on the canonical view garment $C$, the error map $M$, and the pose representation $P$. An overview is shown in Fig. 3, where a Weakly-Supervised Classifier (WSC) first locates the errors in the error map $M$, then $M$ reweights the noise schedule of the I2SB stochastic process and assigns a higher volume of noise to the highlighted low-quality region ”rebel“ such that the model can focus on refining it.
@@
where $\mu_{t}$ is the same as in Eq. 2 and $\epsilon^{r}$ is the adaptive noise.
@@
The initial image $x_{1}$ is iteratively refined to $x_{0}$ via a denoising/sampling process, where a model predicts the noise distribution at each time step. Different from prior soft-attention-based UNets [38, 37, 19, 27], our denoising model uses cloth-flow-learning UNet for more precise garment deformation [15]. It accepts the garment $C$, the error map $M$, the pose representation $P$, and the noisy image $x_{t}$ as inputs and predicts the error-adapted noise $\epsilon^{r}_{\theta}(\cdot;t)$, where $(\cdot;t)$ omits the inputs $M,P,x_{t},C$. See Supp. B for the detailed model architecture. With the predicted noise $\epsilon^{r}_{\theta}(\cdot;t)$, we define our sampling process:
@@ 5.2.1 Further Improvements via Classifier Guidance and Expert Denoisers
Different from prior work that uses object category classifier to guide the sampling process [11], our WSC guidance gives a direction toward the real data distribution. [8] shows that we can estimate the guidance score $\nabla_{x_{t}}\log{p(y|x_{t})}$ using the denoised clean image $\hat{x}_{0}$: $\nabla_{x_{t}}\log{p(y|x_{t})}\simeq\nabla_{x_{t}}\log{p(y|\hat{x}_{0})}$, where $y$ is the fake/real label. Since the label for real data is 0 in WSC, the classifier guidance gives us
@@
SD-Based
@@
Stable-VTON [19] 8.20 0.07 5.16 0.05 0.917 0.057 -
EARSB 8.42 0.07 5.25 0.05 0.918 0.059 10.89 0.13 7.15 0.13 0.961 0.028
EARSB +H2G-UH/FH 8.26 0.06 5.14 0.04 0.919 0.058 10.70 0.11 7.05 0.11 0.965 0.026
EARSB(SD) +H2G-UH/FH 8.04 0.06 4.90 0.03 0.925 0.053 10.41 0.09 6.76 0.08 0.968 0.023
Table 1: Results on VITON-HD [21] and DressCode-Upper [26] KID is multiplied by 100. Results of the diffusion methods are reported at 25 sampling steps.
Further improvements are made via expert denoisers [2], wherein the trained EARSB model is then split into two models, respectively fine-tuned on denoising ranges $t\in[0,0.5]$ and $t\in[0.5,1]$. We default to this in all variants of our method in the next section.
@@ 6 Experiments
We use VITON-HD, DressCode-Upper, and our synthetic H2G-UH and H2G-FH for training. They include 11,647, 13,564, 12,730, 8,939 training images, respectively. For synthetic data augmentation, we combine VITON-HD [21] with our H2G-UH since both of them include mostly upper-body human images. DressCode-Upper is combined with H2G-FH as both consist of full-body human photos. For evaluation, VITON-HD contains 2,032 (human, garment) test pairs and DressCode-Upper has 1,800 test pairs. Experiments are conducted for both paired and unpaired settings. In the paired setting, the input garment image and the garment in the human image are the same item. Conversely, the unpaired setting uses a different garment image.
@@
We use Structural Similarity Index Measure (SSIM) [33], Frechet Inception Distance (FID) [16], Kernel Inception Distance (KID) [4], and Learned Perceptual Image Patch Similarity (LPIPS) [39] to evaluate image quality. All the compared methods use the same image size 512x512 and padding when computing the above metrics.
6.1 Comparison with Existing Methods
We compare our EARSB with GAN-based methods HR-VTON [21], SD-VTON[30] and GP-VTON[36], as well as Stable Diffusion (SD) [29] based methods including CAT-DM [38], Stable-VTON [19], TPD [37], IDM-VTON [6] and CatVTON [7]. We control the number of sampling steps by fixing it to 25 in Tab. 1 for all the diffusion methods. The results under different sampling steps on VITON-HD are presented in Fig. 4.
In Tab. 1, EARSB is our model using GAN method GP-VTON [36] to generate the initial image and is trained without synthetic data augmentation. EARSB+H2G-UH/FH trains with either H2G-UH or H2G-FH. We add the upper-body synthetic subset H2G-UH for the upper-body-human dataset VITON-HD, and the full-body synthetic H2G-FH when on DressCode-Upper. EARSB(SD)+H2G-UH/FH further uses diffusion model CatVTON [7] to generate the initial image. EARSB and EARSB+H2G-UH/FH improve the GAN baselines while EARSB(SD)+H2G-UH/FH gains over the SD-based methods. Note that while using the diffusion model to generate the initial image gives better performance in EARSB(SD)+H2G-UH/FH, it also introduces higher computational costs. Overall, our refinement model improves the metrics of existing methods on both datasets, with further improvements from incorporating synthetic training pairs.
Methods GP-GAN EARSB Stable-VTON EARSB

### v3

@@ Abstract
Given an isolated garment image in a canonical product view and a separate image of a person, the virtual try-on task aims to generate a new image of the person wearing the target garment. Prior virtual try-on works face two major challenges in achieving this goal: a) the paired (human, garment) training data has limited availability; b) generating textures on the human that perfectly match that of the prompted garment is difficult, often resulting in distorted text and faded textures. Our work explores ways to tackle these issues through both synthetic data as well as model refinement. We introduce a garment extraction model that generates (human, synthetic garment) pairs from a single image of a clothed individual. The synthetic pairs can then be used to augment the training of virtual try-on. We also propose an Error-Aware Refinement-based Schrödinger Bridge (EARSB) that surgically targets localized generation errors for correcting the output of a base virtual try-on model. To identify likely errors, we propose a weakly-supervised error classifier that localizes regions for refinement, subsequently augmenting the Schrödinger Bridge’s noise schedule with its confidence heatmap. Experiments on VITON-HD and DressCode-Upper demonstrate that our synthetic data augmentation enhances the performance of prior work, while EARSB improves the overall image quality. In user studies, our model is preferred by the users in an average of 59% of cases. Code is available at this link.
@@ 1 Introduction
Virtual try-on aims to generate a photorealistic image of a target person wearing a prompted product-view garment [23, 41, 38]. It allows users to visualize how garments would fit and appear on their bodies without the need for physical trials. While recent methods have made significant strides in this field [31, 37, 38, 19], noticeable artifacts such as text distortion and faded textures persist in generated images. For example, as illustrated in the second row of Fig. 1, the logo and the text on the t-shirt noticeably fade away in the initial image generated by a prior try-on model [31]. These imperfections stem from two primary challenges in virtual try-on: limited data availability and the complexity of accurate garment texture deformation. To address these issues, we propose a two-pronged approach: augmenting training data through cost-effective synthetic data generation, and surgically targeting known generation artifacts using our proposed Error-Aware Refinement-based Schrödinger Bridge (EARSB).
At a minimum, the training data of virtual try-on requires paired (human, product-view garment) images. The product-view garment image is a canonical, front-facing view of the clothing with a clean background. A substantial amount of data is needed to capture the combinatorial space comprising all possible human poses, skin tones, viewing angles, and their respective physical interactions with fabric textures, shapes, letterings, and other material properties. Unfortunately, these images are generally available only on copyright-protected product webpages and, therefore, are not readily available for use. To mitigate this issue, we propose to augment training with synthetic data generated from the easier symmetric human-to-garment task, wherein we train a garment-extraction model to extract a canonical product-view garment image from an image of a clothed person. This will allow us to create synthetic paired training data from unpaired datasets [25, 12, 36]. Our results demonstrate that incorporating the more readily available synthetic training pairs can improve image generation quality in the virtual try-on task.
@@
Two components are necessary to achieve such a pipeline: a classifier for identifying localized generation errors, and a refinement model that can re-synthesize content specifically in these localized regions. We found that an effective Weakly-Supervised error Classifier (WSC) can be constructed with just a few hours of manual labeling of generation errors. Another benefit of this approach is that it can be easily tailored for the errors of a specific model that produces images with artifacts. The resulting WSC will produce an error map highlighting low-quality regions. Subsequently, we adopt an Image-to-Image Schrodinger Bridge (I2SB) [24] to learn the refinement of these regions in the generated images. While typical diffusion models map from noise to data, I2SB constructs a Schrödinger Bridge (SB) that allows us to map from data to data, or in our setup, generations with artifacts to ground truth images. In addition, we introduce an adaptive noise schedule to direct the SB process to focus on the localized errors by incorporating the classifier’s prediction error into the noise schedule, which we describe in more detail in Sec. 4.1. As shown in the first row of Fig. 1, our refinement SB model (i.e., EARSB) corrects the distorted text in the initially generated image.
@@ 2 Related Work
The addition of synthetic data is often an effective means of improving downstream task performance when it is difficult to amass real data at the necessary scale. This has been demonstrated in the domains of image generation [18, 32] and image editing [5, 35]. Careful applications can also be used to ameliorate dataset imbalance issues, as shown in [10]. Other works, such as [1], use self-synthesized data to provide negative guidance for the diffusion model. Our incorporation of synthetic data in the virtual try-on task tackles a specific sub-problem in the broader image editing domain and is similar in spirit to [5, 35]. Specifically, we aim to synthesize paired training data that satisfies the stringent requirements of virtual try-on paired training data – a canonical product-view garment image paired with an example of it being worn. Images of people in clothing are readily available, but it is difficult to obtain a product-view image of the exact clothing they are wearing. To address this, our work tackles the human-to-garment problem, which aims to extract the clothing from a person’s photo and project it to the canonical product view, making it roughly symmetric to the virtual try-on task.
@@
3 Augmented Training with Synthetic Data
Virtual Try-On Task Definition
@@ 3 Augmented Training with Synthetic Data
Acquiring high-quality pairs $(x_{0},C)$ at scale is challenging due to copyright and brand protection, but acquiring images of just humans ($x_{0}$) at scale is considerably more feasible [25, 12, 36, 22]. This observation motivates proposed human-to-garment process to extract a synthetic canonical view image $\hat{C}$ from $x_{0}$. We can then augment our virtual try-on training with $(x_{0},\hat{C})$ pair, requiring only single human images. In the following, Sec. 3.1 explains the architecture of our human-to-garment model, Sec. 3.2 discusses how we use this model to construct the synthetic dataset, and Sec. 3.3 describes how the synthetic data is used to augment the virtual try-on training.
@@ 3.3 Augmented Virtual Try-on Training
To further prevent distribution leakage of incorporating synthetic data, we explore two means of limiting the effect of the real-synthetic domain gap: (a) two training stages involving pretraining the try-on model using synthetic pairs, and then finetuning on real pairs [20]; (b) training simultaneously on real and synthetic data, but conditioning the try-on model on a real/synthetic flag, similar to [17]. We found empirically that the second augmentation performs slightly better than the first (See Sec. 5.1).
@@ 4 Error-Aware Refinement Schrödinger Bridge
Apart from the synthetic data augmentation from Sec. 3, our second approach to enhancing existing try-on methods is a refinement pipeline with two steps. First, given some base model $F_{base}(\bar{x}_{0},C,\phi)\to x_{1}$ where $F_{base}()$ can be any pretrained GAN or diffusion-based approach to virtual try-on, $x_{1}$ closely approximates the true real human image $x_{0}$ with some generation artifacts. To automatically identify the artifacts in $x_{1}$, we construct a weakly-supervised error classifier $\text{WSC}(x_{1},C)\to M$ as in Fig. 3, where $M$ is a confidence map predicting a heatmap for likely generation errors. Then, our second step performs the final refinement by fitting a Schrödinger bridge based on I2SB [24] via the following mapping: $F_{EARSB}(x_{1},C,M,\phi;\theta)\to x_{0}$.
The approach is weakly inspired by boosting methods in that we wish to fit a targeted refinement model that is trained specifically on the generation errors of an existing model. The refinement goal applies to the general setting where we want to refine a flawed image output, though we focus on virtual try-on for this work. Thus, as illustrated in Fig. 3, the training of EARSB includes three steps:
@@
As the first step simply requires running an off-the-shelf virtual try-on model to obtain an initial image, our discussion in the next section will begin by describing the second step- obtaining the error map.
@@ 4.1 Obtaining the Error Map
We start by obtaining the error map $M$ that highlights the corrupted or incorrect area of the initial image $x_{1}$ using our proposed Weakly-Supervised error Classifier (WSC).
@@
The trained WSC will predict an error map $M$ for the initial image $x_{1}$, which is then used to adjust the noise schedule in the diffusion process described in the next section.
@@ 4.2 Error-Map-Reweighted SB Formulation
Our choice of locally scaling the noise schedule is based on several observations. We want the model to directly copy pixels over to $x_{0}$ for correctly generated regions in $x_{1}$ – it would be nice to avoid training the model to add and remove noise from these regions. In contrast, erroneous regions in the initial try-on images, especially more noticeable ones, include generation errors that share little to no structural similarity to the target. These errors include examples such as deformed limbs, and distorted textures/fabrics, which may need more added noise to prevent the model from conditioning too strongly on the original pattern.
As such, we construct our refinement model $F_{EARSB}(x_{1},C,M,P;\theta)\to x_{0}$, where the model is conditioned on the canonical view garment $C$, the error map $M$, and the pose representation $P$. Fig. 3 shows our Weakly-Supervised Classifier (WSC) first locates the errors in the error map $M$, then $M$ reweights the noise schedule of the I2SB stochastic process to assign a higher volume of noise to the low-quality region “rebel" so the model can focus on refining it.
@@
where $\mu_{t}$ is the same as Eq. 2 and $\epsilon^{r}$ is the adaptive noise.
@@
The initial image $x_{1}$ is iteratively refined to $x_{0}$ via a denoising/sampling process, where a model predicts the noise distribution at each time step. In contrast to prior soft-attention-based UNets [39, 38, 19, 27], our denoising model uses cloth-flow-learning UNet for more precise garment deformation [15]. It accepts the garment $C$, the error map $M$, the pose representation $P$, and the noisy image $x_{t}$ as inputs and predicts the error-adapted noise $\epsilon^{r}_{\theta}(\cdot;t)$, where $(\cdot;t)$ omits the inputs $M,P,x_{t},C$. See Supp. B for the detailed model architecture. With the predicted noise $\epsilon^{r}_{\theta}(\cdot;t)$, we define our sampling process:
@@ 4.2.1 Further Improvements via Classifier Guidance and Expert Denoisers
Whereas prior work used an object category classifier to guide the sampling process [11], our WSC guidance gives a direction toward the real data distribution. Chung et al. [8] shows we can estimate the guidance score $\nabla_{x_{t}}\log{p(y|x_{t})}$ using the denoised clean image $\hat{x}_{0}$: $\nabla_{x_{t}}\log{p(y|x_{t})}\simeq\nabla_{x_{t}}\log{p(y|\hat{x}_{0})}$, where $y$ is the fake/real label. Since the label for real data is 0 in WSC, the classifier guidance is:
@@
EARSB (ours) 8.42 0.07 5.25 0.05 0.918 0.059 10.89 0.13 7.15 0.13 0.961 0.028
EARSB +H2G-UH/FH (ours) 8.26 0.06 5.14 0.04 0.919 0.058 10.70 0.11 7.05 0.11 0.965 0.026
(b) SD-Based
@@
StableVITON [19] 8.20 0.07 5.16 0.05 0.917 0.057 -
EARSB(SD) +H2G-UH/FH (ours) 8.04 0.06 4.90 0.03 0.925 0.053 10.41 0.09 6.76 0.08 0.968 0.023
Table 1: Virtual try-on results on VITON-HD [21] and DressCode-Upper [26] for (a) GAN-based and (b) diffusion-based models using 25 sampling steps. KID is multiplied by 100. We find our EARSB approach outperforms prior work on average. See Sec. 5.1 for discussion.
Following [2], a trained EARSB model is split into two models, each having an expert denoiser that is fine-tuned on denoising ranges $t\in[0,0.5]$ and $t\in[0.5,1]$, respectively.
@@ 5 Experiments
We use VITON-HD [21], DressCode-Upper [26], and our synthetic H2G-UH and H2G-FH for training. They include 11,647, 13,564, 12,730, 8,939 training images, respectively. For synthetic data augmentation, we combine VITON-HD with our H2G-UH since they both include mostly upper-body human images. DressCode-Upper is combined with H2G-FH as both consist of full-body human photos. For evaluation, VITON-HD contains 2,032 (human, garment) test pairs and DressCode-Upper has 1,800 test pairs and include both paired and unpaired settings. In the paired setting, the input garment image and the garment in the human image are the same item. Conversely, the unpaired setting uses a different garment image.
@@
We use Structural Similarity Index Measure (SSIM) [34], Frechet Inception Distance (FID) [16], Kernel Inception Distance (KID) [4], and Learned Perceptual Image Patch Similarity (LPIPS) [40] to evaluate image quality. All the compared methods use the same image size $512\times 512$ and padding when computing the above metrics.
Experimental setup
We compare EARSB with GAN-based methods HR-VTON [21], SD-VTON [31] and GP-VTON [37], as well as Stable Diffusion (SD) based methods including CAT-DM [39], StableVITON [19], TPD [38], IDM-VTON [6] and CatVTON [7]. Unless otherwise specified, all diffusion models use 25 sampling steps.
We report results of multiple variants of our approach. EARSB uses GAN-based GP-VTON [37] to generate the initial image that was trained without synthetic data augmentation. EARSB+H2G-UH/FH trains with either H2G-UH or H2G-FH. We add the upper-body synthetic subset H2G-UH for the upper-body-human dataset VITON-HD, and the full-body synthetic H2G-FH when on DressCode-Upper. EARSB(SD)+H2G-UH/FH uses the diffusion-based CatVTON [7] to generate the initial image.
5.1 Results
Tab. 1 compares our approach with those from prior work. We find our full model variants EARSB+H2G-UH/FH and EARSB(SD)+H2G-UH/FH boosts performance over the GAN and SD-based methods, respectively. The last two rows of Tab. 1(a) show that incorporating our synthetic training pairs provide a consistent boost in performance on both datasets. Comparing the last rows of Tab. 1(a) and Tab. 1(b) we observe that using the diffusion model to generate the initial image gives better performance in EARSB(SD)+H2G-UH/FH, but is more costly.
Methods GP-VTON EARSB StableVITON EARSB